OpenStudy (anonymous):

please help me | | v v Identify the solutions of 3x2 − 7x + 4 = 0. x = −1 over 3, x = −4 x = 1 over 3, x = 4 x = −4 over 3, x = −1 x = 4 over 3, x = 1

OpenStudy (anonymous):

(3x-4)(x-1) x = 4/3 or x=1

OpenStudy (anonymous):

thank you so much can you help me solve other one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

on the same on or should i close this one?

OpenStudy (anonymous):

this is fine

OpenStudy (anonymous):

ok, ] Using graphing technology, approximate the solutions of 0.91x2 − 0.54x − 4.9 = 0. x ≈ −2.32 and x ≈ 2.56 x ≈ −2.14 and x ≈ 2.44 x ≈ −2.18 and x ≈ 2.53 x ≈ −2.04 and x ≈ 2.64

OpenStudy (anonymous):

give me a minute

OpenStudy (anonymous):

yes ma' am

OpenStudy (anonymous):

OK...I cheated and used a quadratic formula calculator that I found online. The answer is the 4th choice.

OpenStudy (anonymous):

ok thank you if this is too much let me know but i have 9 more questions is it ok if you help me do does if you do i will be so greatful and happy please and thank you (but it is your choice)

OpenStudy (anonymous):

I have a few more minutes

OpenStudy (anonymous):

ok thank you i will give them to you all at once and you can just give me the answer all at once if it is ok with you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What are the approximate solutions of 2x2 + 5x = 4 rounded to the nearest hundredth? No real solutions x ≈ −6.27 and x ≈ 1.27 x ≈ −0.64 and x ≈ 3.14 x ≈ −3.14 and x ≈ 0.64 Question 14 (Multiple Choice Worth 3 points) [9.02] What are the solutions of 4x2 = −20x? x = 0, x = 5 x = 0, x = −5 x = 4, x = 5 x = 4, x = −5 Question 15 (Multiple Choice Worth 4 points) [9.01] Determine whether the graph of y = x2 − 4x + 1 has a maximum or minimum point, then find the maximum or minimum value. Minimum; (2, -3) Maximum; (2, -3) Minimum; (-3, 2) Maximum; (-3, 2) Question 16 (Multiple Choice Worth 4 points) [9.06] A firecracker shoots up from a hill 165 feet high with an initial speed of 110 feet per second. Using the formula H(t) = −16t2 + vt + s, determine how long it will take the firecracker to hit the ground. 7.2 seconds 7.5 seconds 7.8 seconds 8.1 seconds Question 17 (Multiple Choice Worth 3 points) [9.05] How many and of what type are the solutions of a quadratic equation when the value of the radicand is 0? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions Question 18 (Multiple Choice Worth 4 points) [9.02] Where does the graph of y = 2x2 − 5x + 3 cross the x-axis? (−1 over 2, 0) and (−3, 0) (1 over 2, 0) and (3, 0) (3 over 2, 0) and (1, 0) (−3 over 2, 0) and (−1, 0) Question 19 (Multiple Choice Worth 3 points) [9.01] Which equation is a quadratic equation? y + 7 = −3x(4 − x) + 9x y = −2x + 5 y + 2x3 = (2x − 2)(x + 5) y − 9x = (x2 + 1)(5x − 1) + 3 Question 20 (Multiple Choice Worth 3 points) [9.06] Katie wants to hang a painting in a gallery but the painting and frame must have an area of 45 square feet. The painting is 6 feet wide by 7 feet long. The frame must have an equal width of x on each side. Which quadratic equation represents the area of the painting and frame combined? 4x2 + 26x + 45 = 0 4x2 + 26x − 3 = 0 x2 + 13x − 3 = 0 x2 + 13x + 45 = 0 Question 21 (Multiple Choice Worth 3 points) [9.01] Determine whether the graph of y = −2x2 + 7x − 10 opens up or down and whether it has a maximum or minimum point. Up; Maximum Up; Minimum Down; Maximum Down; Minimum Question 22 (Multiple Choice Worth 4 points) [9.01] Identify the vertex for the graph of y = −2x2 − 12x + 5. (3, −13) (−3, −13) (−3, 23) (3, 23) Question 23 (Multiple Choice Worth 4 points) [9.05] What are the exact solutions of x2 + 5x + 2 = 0? x = the quantity of negative 5 plus or minus the square root of 33 all over 2 x = the quantity of 5 plus or minus the square root of 33 all over 2 x = the quantity of negative 5 plus or minus the square root of 17 all over 2 x = the quantity of 5 plus or minus the square root of 17 all over 2

OpenStudy (anonymous):

the first one is x=0.64 or x = -3.14 (last choice)

OpenStudy (anonymous):

9.02 x = 0 or x = -5 second choice

OpenStudy (anonymous):

9.01 minimum (2,-3)

OpenStudy (anonymous):

thank you so much are you still working on the rest ma'am?

OpenStudy (anonymous):

Don't know about the velocity problem

OpenStudy (anonymous):

9.05 is the second answer --2 identical rational solutions

OpenStudy (anonymous):

9.02 (3/2,0)(1,0) 3rd one

OpenStudy (anonymous):

thank you so far so good!!!

OpenStudy (anonymous):

9.01 not sure--I'd guess the 2nd one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

9.01 maximium and down

OpenStudy (anonymous):

oh actually for the one you guessed on is the first one thnx anyways

OpenStudy (anonymous):

9.01 (-3,23)

OpenStudy (anonymous):

the last problem--3rd choice

OpenStudy (anonymous):

thank you so much ma'am

OpenStudy (anonymous):

you're welcome. Good luck